OpenStudy (anonymous):

Each side of a square is increasing at a rate of 6 cm/s. At what rate is the area of the square increasing when the area of the square is 16 cm^2?

OpenStudy (anonymous):

\[A_\square = s^2\] where s is the length of a side. \[ds/dt = 6 cm/s\] take the derivitive of the area \[dA/dt = 2 *s *ds/dt\] we are given that area is 16, so the length of a side (s) is 4 \[dA/dt = 2 *4 *6\]